OpenStudy (alibaby):

There are (10^8)^2 ⋅ 10^0 candies in a store. What is the total number of candies in the store? 1 10^10 10^16 10^17

OpenStudy (alibaby):

i think it's C

OpenStudy (alibaby):

but the answer to this problem is 1x10^16

OpenStudy (alibaby):

@Will.H

OpenStudy (leenathan):

not will

OpenStudy (alibaby):

so i dont know which one is it

OpenStudy (will.h):

you are correct it is C Do you need explanation or do you know why ?

OpenStudy (phi):

1x10^16 is the same as 10^16 just like 1*2 is the same as 2

OpenStudy (mathstudent55):

Must the answer be written in scientific notation?

OpenStudy (alibaby):

id like an explanation plz

OpenStudy (phi):

\[ \left( 10^8\right)^2 = (10^8) (10^8) = 10^{16} \]

OpenStudy (mathstudent55):

\(10^{16} = 1 \times 10^{16}\) \(10^{16} \) is not written in scientific notation. \(1 \times 10^{16} \) is written in scientific notation.

OpenStudy (will.h):

\[(10^8)^2 * 10^0\] Simplify the left side How? multiply the exponents 8^2 = 16 and the right side basically equals to 1

OpenStudy (phi):

and of course 10^0 is 1 if the answer key says 1*10^16 but the choices leave off the 1, then they are being sloppy

OpenStudy (will.h):

Lol yeah

OpenStudy (mathstudent55):

A number written in scientific notation consists of the product of a number from 1 to under 10 and an integer power of 10.

OpenStudy (alibaby):

hehe ummmm ok lol thanks @phi @Will.H @mathstudent55

OpenStudy (will.h):

you welcome